OpenStudy (anonymous):

The surface area of a spherical balloon is 175 cm2. If the length of the radius was six times as large, what would the surface area be?

OpenStudy (shayaan_mustafa):

first calculate radius from\[\large A=4\pi r^2\]

OpenStudy (anonymous):

how would i get the radius?

OpenStudy (anonymous):

surface area= 4*pi*r^2 radius would be 6*r, so in our formula r^2, thus (6r)^2=36r^2, so multiply the given surface area by 36

OpenStudy (anonymous):

so my answer should be 6300?

OpenStudy (shayaan_mustafa):

you have given with A=175 put it in the formula I have given.\[\large 175=4\pi r^2\]\[\large \frac{175}{4\pi}=r^2\]

OpenStudy (shayaan_mustafa):

calculate r from here.

OpenStudy (shayaan_mustafa):

by calculation you would get \[\large r=\pm3.73cm\]but radius is the length. So length can never be negative. choose the +\[\large r=3.73cm\]Now you have got your radius.

OpenStudy (shayaan_mustafa):

Is it clear up till now?

OpenStudy (shayaan_mustafa):

@kkaila1

OpenStudy (shayaan_mustafa):

I have to go kkaila1 Bye. Tc :)